# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ 4 Accelerated convergence under strong convexity
Under Assumption 2.4, further assume that $D_{w^{x}}(x,x^{\prime})\leq\frac{C}{2}\|x-x^{\prime}\|^{2}$ for all $x,x^{\prime}\in\mathcal{X}$ for some $1\leq C<\infty$, and $\mu>0$. Let $\{\rho_{s}\}_{s\in[S]}$ be a distribution over $[S]$, $\overline{r}=\sum_{s=1}^{S}r_{s}\rho_{s}$ and similarly define $\overline{r^{2}}$ and $\overline{r^{3}}$.
With $\alpha_{s,i_{s}}=1$, $\theta_{k}=k+2\overline{r^{2}}/\overline{r}$, $\eta_{k}=\frac{\mu}{2\overline{r}C}(k+\overline{r^{2}}/\overline{r})$, $\eta_{k,s}=\eta_{k}\rho_{s}$, $\tau_{s,i_{s}}(\sum_{k^{\prime}=r_{s}i_{s}}^{r_{s}i_{s}+r_{s}-1}\theta_{k^{\prime}})=\tau_{s}=\frac{\widetilde{\kappa}_{s}^{2}}{\rho_{s}}\cdot\frac{4r_{s}\overline{r^{2}}C}{\mu}$, $T_{k}/N=T/N\geq\max(\frac{5}{\sqrt{D_{1}}},\frac{64\overline{r}}{D_{1}})$ where $D_{1}=\frac{\mu^{2}(\overline{r^{2}}/\overline{r})^{2}}{2M^{2}C^{2}}D_{0}$, assume that the CS procedure at iteration $k$ is run with $\lambda_{t}=t$ and $\beta^{k}_{t}=\frac{(t+1)\mu}{2\eta_{k}C}+\frac{t-1}{2}$ for $t=1,\ldots,T_{k}$, then for any $Z\in\overline{\mathcal{Z}}$
[…]
$+\frac{m\mu(\overline{r^{2}}/\overline{r})^{2}}{C}D_{0}+\frac{4\overline{r}C}{\mu}\sum_{s=1}^{S}\frac{\widetilde{\kappa}_{s}^{2}r_{s}}{\rho_{s}}D_{w^{y}_{s}}(y_{s},y_{s}^{init})\biggr{\}}.$
As a direct consequence, we have the following theorem.
Theorem 4.1.
$D_{w^{X}}(\widehat{X},X^{init})\leq D^{X}<\infty,\quad\sup_{y_{s}\in\mathsf{dom}(R^{*}_{s})}D_{w^{y}_{s}}(y_{s},y_{s}^{init})\leq D_{s}^{y}<\infty.$
Under the conditions in Lemma 4.1, taking $\rho_{s}=\frac{\widetilde{\kappa}_{s}\sqrt{D_{s}^{y}}}{\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}}}$ and $D_{0}=D^{X}/m$,
$\sup_{Y^{\prime}\in\mathbb{R}^{n}}Q(Z^{N};\widehat{X},Y^{\prime})\leq\frac{2}{N(N+1)}\left\{\frac{\mu(\overline{r^{3}}/\overline{r}+7(\overline{r^{2}}/\overline{r})^{2})}{2C}D^{X}+\frac{4C(\overline{r})^{2}}{\mu}(\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}})^{2}\right\}.$
Notice that (1) implies that $\|x-x^{\prime}\|\leq\frac{M}{\mu}$ for all $x,x^{\prime}\in\mathcal{X}$. Thus, one can take $D^{X}=O(\frac{mCM^{2}}{\mu^{2}})$. The resulting upper bound, when $\overline{r^{3}}=O((\overline{r})^{3})$ and $\overline{r^{2}}=O((\overline{r})^{2})$, becomes
$\sup_{Y^{\prime}\in\mathbb{R}^{n}}Q(Z^{N};\widehat{X},Y^{\prime})=O(\frac{\overline{r}^{2}}{\mu N^{2}}\left\{mM^{2}+C(\sum_{s^{\prime}=1}^{S}\widetilde{\kappa}_{s^{\prime}}\sqrt{D_{s^{\prime}}^{y}})^{2}\right\}).$
4.1 Good initialization for ($\mathcal{P}$)
In Theorem 4.1, assuming that $\mathcal{X}$ is compact, then one can always use $D^{X}\geq\sup_{X\in\overline{\mathcal{X}}}D_{w^{X}}(X,X^{init})$, suggesting that $X^{init}$ should be chosen as the “center” of $\overline{\mathcal{X}}$, and $D^{X}$ measures the (squared) radius of $\overline{\mathcal{X}}$. The resulting $N$, then, depends on $D^{X}$. However, such dependence on the size of $\overline{\mathcal{X}}$ could be suboptimal, especially when local objectives are similar. Indeed, in the extreme case where all local functions are the same, then primal agents can optimize their local objectives without communication at all.
To take advantage of potential similarities in the local functions, we propose initializing the primal variables at (approximate) local optimal solutions, which has the following guarantee on $D_{w^{X}}(X^{*},X^{init})$.
Lemma 4.2.
Assume that all norms are the $l_{2}$ norm, and for some $\epsilon_{0}\geq 0$, $\widehat{X}=(\widehat{x}_{v})_{v\in V}\in\overline{\mathcal{X}}$ satisfies the following condition
$F(\widehat{X})\leq\min_{X\in\overline{\mathcal{X}}}F(X)+\epsilon_{0}.$
Assume that (1) holds for some $\mu>0$ and Assumption 2.1 holds, and $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar, then
@@ 4.1 Good initialization for ($\mathcal{P}$)
The above Lemma 4.2 shows that if $\{\widehat{x}_{v}\}_{v\in V}$ are all approximately optimal to local objectives, then $D_{w^{X}}(X^{*},\widehat{X})\sim\frac{\sum_{s=1}^{S}a_{s}^{2}}{\mu^{2}}$. To find such initialization, one can apply the CS procedure.
@@
Assume that all norms are $l_{2}$ norms and $w^{x}(x)=\frac{1}{2}\|x\|^{2}$, that (1) holds with some $\mu>0$, and that Assumption 2.1 holds. For each $v\in V$, assume that $\mathsf{Agent}(x_{v})$ is given some $\underline{x}_{v}^{0}\in\mathcal{X}$ such that $\sup_{x\in\mathcal{X}}D_{w^{x}}(x,\underline{x}_{v}^{0})\leq\underline{D}^{x}<\infty$
(_$(\_,x_{v}^{init})=CS(f_{v},\mathcal{X},D_{w^{x}},\underline{T},\underline{\eta},\mathbf{0},\underline{x}_{v}^{0},\underline{x}_{v}^{0}),$
where the CS procedure uses $\lambda_{t},\beta_{t}$ according to Corollary 3.1 (for $\mu>0$), then with $\epsilon_{0}=\widetilde{a}^{2}/\mu$, $\underline{T}\geq\frac{8CM^{2}m}{\epsilon_{0}\mu}$ and $\underline{\eta}=\frac{\epsilon_{0}/2}{m\underline{D}^{x}}$
$D_{w^{X}}(X^{*},X^{init})\leq\frac{4\widetilde{a}^{2}}{\mu^{2}},$
@@
4.2 Complexities for ($\mathcal{P}$) and discussions
Recall that with $w_{s}^{y}=\frac{1}{2}\|y_{s}\|_{2}^{2}$ and $y_{s}^{0}=\mathbf{0}$, we can take $\sqrt{2D_{s}^{y}}$ as $\frac{\xi+a_{s}}{\sigma_{\min}^{+}(K_{s})}$ for ${R}_{s}^{ccv}$ in (5), as $\frac{(1+\xi)a_{s}}{\sigma_{\min}^{+}(K_{s})}$ for ${R}_{s}^{prj}$ in (6), as $\frac{\xi+\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$ for $\widehat{R}_{s}^{ccv}$ in (12), and as $\frac{(1+\xi)\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$ for $\widehat{R}_{s}^{prj}$ in (13). Now, combining Theorem 4.1, Corollary 4.1, Lemma 2.3, and Corollary 2.1, we get the following results.
@@ 4.2 Complexities for ($\mathcal{P}$) and discussions
Assume that all norms are the $l_{2}$ norm, and take $y_{s}^{0}=\mathbf{0}$, $w_{s}^{y}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$ and $w^{x}(x)=\frac{1}{2}\|x\|^{2}$. Assume that Assumptions 2.1, 2.2 and 2.3 and the conditions of Theorem 4.1 and Corollary 4.1 hold, and $\{f_{v}\}_{v\in V}$ is $\widehat{a}_{1}$-similar for some $\widehat{a}_{1}>0$, and take $\rho_{s}=\frac{\|K_{s}\|}{\sum_{s^{\prime}=1}^{S}\|K_{s^{\prime}}\|}$. In addition, assume that $X^{init}$ is initialized according to Corollary 4.1, then we have
$\sup_{Y^{\prime}\in\mathbb{R}^{n}}Q(Z^{N};X^{*},Y^{\prime})\leq\frac{4(\overline{r})^{2}}{\mu N^{2}}\left\{(\overline{r^{3}}/(\overline{r})^{3}+7(\overline{r^{2}}/(\overline{r})^{2})^{2})\cdot\widehat{a}_{1}^{2}+A_{0}^{2}\right\},$
and so for
$N\geq\frac{2\overline{r}A}{\sqrt{\mu\epsilon}},\quad A=\sqrt{\overline{r^{3}}/(\overline{r})^{3}+7(\overline{r^{2}}/(\overline{r})^{2})^{2}}\cdot\widehat{a}_{1}+A_{0},$
@@
    $\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=\widehat{R}_{s}^{ccv}$ as defined in (12) and $A_{0}=\frac{(\sum_{s=1}^{S}\|K_{s}\|)}{\sigma_{\min}^{+}(K)}\cdot(\xi+\widehat{a}_{1})$;
@@
    $\Pi(\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=\widehat{R}_{s}^{prj}$ as defined in (13) and $A_{0}=\frac{(1+\xi)(\sum_{s=1}^{S}\|K_{s}\|)\cdot\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$.
Corollary 4.3.
Assume that all norms are the $l_{2}$ norm, and take $y_{s}^{0}=\mathbf{0}$, $w_{s}^{y}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$ and $w^{x}(x)=\frac{1}{2}\|x\|^{2}$. Assume that $K_{s}K_{s^{\prime}}^{*}=\mathbf{0}$ for all $s,s^{\prime}\in[S]$, that Assumptions 2.1, 2.2 and 2.3 and the conditions of Theorem 4.1 and Corollary 4.1 hold, and that $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar where $a_{s}>0$ for all $s$. In addition, assume that $X^{init}$ is initialized according to Corollary 4.1, then we have
$\sup_{Y^{\prime}\in\mathbb{R}^{n}}Q(Z^{N};X^{*},Y^{\prime})\leq\frac{4(\overline{r})^{2}}{\mu N^{2}}\left\{(\overline{r^{3}}/(\overline{r})^{3}+7(\overline{r^{2}}/(\overline{r})^{2})^{2})\cdot(\sum_{s=1}^{S}a_{s}^{2})+A_{0}^{2}\right\},$
and so for
$N\geq\frac{2\overline{r}A}{\sqrt{\mu\epsilon}},\quad A=\sqrt{\overline{r^{3}}/(\overline{r})^{3}+7(\overline{r^{2}}/(\overline{r})^{2})^{2}}\cdot(\sum_{s=1}^{S}a_{s}^{2})^{1/2}+A_{0},$
  1. 1.
@@
    $\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=R_{s}^{ccv}$ as defined in (5), $\rho_{s}=(\frac{\xi+a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{\xi+a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})^{+}})$, and $A_{0}=\sum_{s=1}^{S}(\xi+a_{s})\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})}$;
  2. 2.
@@
    $\Pi(\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=R_{s}^{prj}$ satisfies (6), $\rho_{s}=(\frac{a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})^{+}})$, and $A_{0}=(1+\xi)(\sum_{s=1}^{S}a_{s}\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})})$.
@@
Subgradient oracle complexities. With the initialization in Corollary 4.1 and $C=1$, the number of subgradient steps needed to find $X^{init}$ is $\underline{T}\geq\frac{8mM^{2}}{\widetilde{a}^{2}}$, constant in $\epsilon$.
@@
In addition, in Theorem 4.1, we can take $D_{0}=\frac{4\widetilde{a}^{2}}{\mu^{2}}$, and so $D_{1}=\frac{2\widetilde{a}^{2}(\overline{r^{2}}/\overline{r})^{2}}{M^{2}}$. Thus, the requirement on $T$ becomes $T/N\geq\max(\frac{5}{\sqrt{D_{1}}},\frac{64\overline{r}}{D_{1}})$, i.e. $T/N=\Omega(\max(\frac{M/\widetilde{a}}{\overline{r^{2}}/\overline{r}},(\frac{M/\widetilde{a}}{\overline{r^{2}}/\overline{r}})^{2}\cdot\overline{r}))$, and so the total subgradient steps needed (for each agent) is
$N^{2}\cdot O(\max(\frac{M/\widetilde{a}}{\overline{r^{2}}/\overline{r}},(\frac{M/\widetilde{a}}{\overline{r^{2}}/\overline{r}})^{2}\cdot\overline{r}))=O(\frac{\overline{r}^{2}A^{2}}{\mu\epsilon}\cdot\max(\frac{M/\widetilde{a}}{\overline{r^{2}}/\overline{r}},(\frac{M/\widetilde{a}}{\overline{r^{2}}/\overline{r}})^{2}\cdot\overline{r}))$
@@
In the special case where $S=1$ and $\|K\|=O(\sigma_{\min}^{+}(K))$, we have $\widetilde{a}=\Omega(A)$. Further assuming that $A=O(\sqrt{m}M)$ (which holds when $\|f^{\prime}_{v}\|\leq M_{f}$ for all $v$ and $M=2M_{f}$), the above can be simplified as $\frac{\overline{r}\sqrt{m}M^{2}}{\mu\epsilon}$.
@@
Communication rounds complexities. For both Corollaries 4.2 and 4.3, the communication rounds needed is $N=O(\frac{\overline{r}A}{\sqrt{\mu\epsilon}})$, where $A$ depends on function similarities and higher moments of $\{r_{s}\}_{s\in[S]}$. In terms of $N$ and $T$’s dependency on $\epsilon,\mu$, $O(1/\sqrt{\mu\epsilon})$ communication rounds and $O(1/(\mu\epsilon))$ gradient steps are needed.
@@
$N_{1}=O\left(\frac{\overline{r}\widehat{a}_{1}}{\sqrt{\epsilon\mu}}\cdot\frac{\sum_{s=1}^{S}\|K_{s}\|}{\sigma_{\min}^{+}(K)}\right),\quad N_{2}=O\left(\frac{\overline{r}}{\sqrt{\epsilon\mu}}\cdot\left(\sum_{s=1}^{S}a_{s}\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})}\right)\right),$

### v2

@@ 4.6 Convergence for strongly convex objectives
Assume that all norms are the $l_{2}$ norm, and for some $\epsilon_{0}\geq 0$, $\widehat{X}=(\widehat{x}_{v})_{v\in V}\in{\mathcal{X}}$ satisfies the following condition
$F(\widehat{X})\leq\min_{X\in{\mathcal{X}}}F(X)+\epsilon_{0}.$
Assume that (17) holds for some $\mu>0$ and $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar, then
@@
The above Lemma 4.4 shows that if $\{\widehat{x}_{v}\}_{v\in V}$ are all approximately optimal to local objectives, then $D(X^{*},\widehat{X})\sim\frac{\sum_{s=1}^{S}a_{s}^{2}}{\mu^{2}}$. To find such initialization, one can apply the GS procedure.
@@
Assume that all norms are $l_{2}$ norms and $w_{x}(x)=\frac{1}{2}\|x\|^{2}$, that (17) holds with some $\mu>0$. For each $v\in V$, assume that $\mathsf{Agent}(x_{v})$ is given some $\underline{x}_{v}^{0}\in\mathcal{X}$ such that $\sup_{x\in\mathcal{X}}D(x,\underline{x}_{v}^{0})\leq\underline{D}^{x}<\infty$
$(\_,x_{v}^{init})=GS(f_{v},\mathcal{X},D,\underline{T},\underline{\eta},\mathbf{0},\underline{x}_{v}^{0},\underline{x}_{v}^{0}),$
where the GS procedure uses $\lambda_{t},\beta_{t}$ according to Corollary A.1 (for $\mu>0$), then with $\epsilon_{0}=\widetilde{a}^{2}/\mu$, $\underline{T}\geq\frac{8CM_{f}^{2}m}{\epsilon_{0}\mu}$ and $\underline{\eta}=\frac{\epsilon_{0}/2}{m\underline{D}^{x}}$
$D(X^{*},X^{init})\leq\frac{4\widetilde{a}^{2}}{\mu^{2}},$
@@
Complexities for ($\mathcal{P}_{d}$). Combining Theorem 3.5, Corollary 4.3, Lemma 4.3, and Corollary 4.1, we get the following results.
@@
Assume that all norms are the $l_{2}$ norm, and take $y_{s}^{0}=\mathbf{0}$, $w_{y_{s}}(y_{s})=\frac{1}{2}\|y_{s}\|^{2}$, $w_{x}(x)=\frac{1}{2}\|x\|^{2}$, and $w_{X}(X)=\sum_{v\in V}w_{x}(x_{v})$. Assume that conditions of Corollaries 3.5 and 4.3 hold and $X^{init}$ is initialized according to Corollary 4.3, and the following holds for $A_{0},A_{1}$ specified below
$N\geq\frac{2\overline{r}A}{\sqrt{\mu_{f}\epsilon}},\quad A=\sqrt{\overline{r^{3}}/(\overline{r})^{3}+7(\overline{r^{2}}/(\overline{r})^{2})^{2}}\cdot A_{1}+A_{0}.$
@@
    If $\{f_{v}\}_{v\in V}$ is $\widehat{a}_{1}$-similar, then take $A_{1}=\widehat{a}_{1}$ and $\rho_{s}=\frac{\|K_{s}\|}{\sum_{s^{\prime}=1}^{S}\|K_{s^{\prime}}\|}$
    1. (a)
      $\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=\widehat{R}_{s}^{ccv}$ as defined in (27) (then $\sqrt{2D_{s}^{y}}=\frac{\xi+\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$) and $A_{0}=\frac{(\sum_{s=1}^{S}\|K_{s}\|)}{\sigma_{\min}^{+}(K)}\cdot(\xi+\widehat{a}_{1})$;
    2. (b)
      $\Pi(\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=\widehat{R}_{s}^{prj}$ as defined in (28) (then $\sqrt{2D_{s}^{y}}=\frac{(1+\xi)\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$) and $A_{0}=\frac{(1+\xi)(\sum_{s=1}^{S}\|K_{s}\|)\cdot\widehat{a}_{1}}{\sigma_{\min}^{+}(K)}$.
@@
    If $\{f_{v}\}_{v\in V}$ is $\{(a_{s},K_{s})\}_{s\in[S]}$-similar where $a_{s}>0$ for all $s$, then take $A_{1}=(\sum_{s=1}^{S}a_{s}^{2})^{1/2}$,
    1. (a)
@@
      $\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}}$ is an $(\epsilon,\epsilon/\xi)$-solution if $R_{s}=R_{s}^{ccv}$ as defined in (20) (then $\sqrt{2D_{s}^{y}}=\frac{\xi+a_{s}}{\sigma_{\min}^{+}(K_{s})}$), $\rho_{s}=(\frac{\xi+a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{\xi+a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})^{+}})$, and $A_{0}=\sum_{s=1}^{S}(\xi+a_{s})\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})}$;
    2. (b)
@@
      $\Pi(\frac{\sum_{k=0}^{N}\theta_{k}\widehat{X}^{k}}{\sum_{k=0}^{N}\theta_{k}})$ is an $(\epsilon(1+1/\xi),0)$-solution if $R_{s}=R_{s}^{prj}$ satisfies (21) (then $\sqrt{2D_{s}^{y}}=\frac{(1+\xi)a_{s}}{\sigma_{\min}^{+}(K_{s})}$), $\rho_{s}=(\frac{a_{s}}{\sigma^{+}_{\min}(K_{s})})/(\sum_{s^{\prime}=1}^{S}\frac{a_{s^{\prime}}}{\sigma^{+}_{\min}(K_{s^{\prime}})^{+}})$, and $A_{0}=(1+\xi)(\sum_{s=1}^{S}a_{s}\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})})$.
@@
Subgradient oracle complexities. With the initialization in Corollary 4.3 and $C=1$, the number of subgradient steps needed to find $X^{init}$ is $\underline{T}\geq\frac{8mM_{f}^{2}}{\widetilde{a}^{2}}$, constant in $\epsilon$.
@@
In addition, in Corollary 3.4, we can take $D_{0}=\frac{4\widetilde{a}^{2}}{\mu_{f}^{2}}$, and so $D_{1}=\frac{2\widetilde{a}^{2}(\overline{r^{2}}/\overline{r})^{2}}{M_{f}^{2}}$. Thus, the requirement on $T$ becomes $T/N\geq\max(\frac{5}{\sqrt{D_{1}}},\frac{64\overline{r}}{D_{1}})$, i.e. $T/N=\Omega(\max(\frac{M_{f}/\widetilde{a}}{\overline{r^{2}}/\overline{r}},(\frac{M_{f}/\widetilde{a}}{\overline{r^{2}}/\overline{r}})^{2}\cdot\overline{r}))$, and so the total subgradient steps needed (for each agent) is
$N^{2}\cdot O(\max(\frac{M_{f}/\widetilde{a}}{\overline{r^{2}}/\overline{r}},(\frac{M_{f}/\widetilde{a}}{\overline{r^{2}}/\overline{r}})^{2}\cdot\overline{r}))=O(\frac{\overline{r}^{2}A^{2}}{\mu_{f}\epsilon}\cdot\max(\frac{M_{f}/\widetilde{a}}{\overline{r^{2}}/\overline{r}},(\frac{M_{f}/\widetilde{a}}{\overline{r^{2}}/\overline{r}})^{2}\cdot\overline{r}))$
@@
In the special case where $S=1$ and $\|K\|=O(\sigma_{\min}^{+}(K))$, we have $\widetilde{a}=\Omega(A)$. Further assuming that $A=O(\sqrt{m}M_{f})$ (which holds when $\|f^{\prime}_{v}\|\leq M_{f}$ for all $v$), the above can be simplified as $\frac{\overline{r}\sqrt{m}M_{f}^{2}}{\mu_{f}\epsilon}$.
@@
Communication rounds complexities. From Corollary 4.4, the communication rounds needed is $N=O(\frac{\overline{r}A}{\sqrt{\mu_{f}\epsilon}})$, where $A$ depends on function similarities and higher moments of $\{r_{s}\}_{s\in[S]}$. In terms of $N$ and $T$’s dependency on $\epsilon,\mu_{f}$, $O(1/\sqrt{\mu_{f}\epsilon})$ communication rounds and $O(1/(\mu_{f}\epsilon))$ gradient steps are needed.
@@
$N_{1}=O\left(\frac{\overline{r}\widehat{a}_{1}}{\sqrt{\epsilon\mu_{f}}}\cdot\frac{\sum_{s=1}^{S}\|K_{s}\|}{\sigma_{\min}^{+}(K)}\right),\quad N_{2}=O\left(\frac{\overline{r}}{\sqrt{\epsilon\mu_{f}}}\cdot\left(\sum_{s=1}^{S}a_{s}\cdot\frac{\|K_{s}\|}{\sigma_{\min}^{+}(K_{s})}\right)\right),$